What is root mean square velocity.


1
Expert's answer
2021-04-23T07:26:51-0400

The term root mean square velocity is applied to the velocity of atoms or molecules in the gas. It is defined as the square root of

"\displaystyle <v^2> = \int_0 ^\infty v^2 f(v) dv"

where f(v) is Maxwell distribution.

After the computation of this internal, one gets formula

"\displaystyle \sqrt{<v^2>} = \sqrt{\frac{3kT}{m}}"

where m is a mass of atom/molecula.